find the radius of the base of a cylinder whose surface area is 603.2 square meters with a height of 10 meters
Answer by stanbon(75887)

find the radius of the base of a cylinder whose surface area is 603.2 square meters with a height of 10 meters
----------------------
Surface are = 2(pi)r*h = 603.2
2pir*10 = 603.2
2pir = 60.32
pir = 30.16
r = 30.16/pi is approx 9.6002 meters
========================================
Cheers,
Stan H.